# Liquidity fragmentation on decentralized exchanges

### v5

@@ 2.2 Pool fragmentation and market quality
In our model, liquidity takers incur two main costs beyond gas prices: pool fees, which represent the cost of taking liquidity, as well as foregone gains from trade if liquidity providers do not fully participate in the market. To integrate these costs into a single measure of market quality, we use an implementation shortfall metric, following Perold (1988). If the asset is traded on a sequence of pools, where $f_{k}$ and $\mathcal{L}_{k}$ represent the fees and liquidity deposits on pool $k$, respectively, the implementation shortfall (IS) for a large LT is defined as
$\text{IS}\left(\left\{f_{k}\right\}_{k}\right)=\underbrace{\sum_{k}f_{k}\mathcal{L}_{k}}_{\text{trading fees}}+\underbrace{\Delta\left(\Theta-\sum_{k}\mathcal{L}_{k}\right)}_{\text{unrealized gains}},$ (13)
where $\Theta-\sum_{k}\mathcal{L}_{k}$ is the difference between the LT trading demand and the cumulative liquidity available across all pools, and $\Delta$ is the per-unit gains from trade.
Suppose an asset is traded on a single pool that imposes a liquidity fee $f$. From equation (13) it follows that the implementation shortfall on this pool is:
$\text{IS}\left(\left\{f\right\}\right)$ $=f\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i+\Delta\left(\Theta-\int_{\frac{\Gamma}{f}}^{Q}q_{i}\varphi\left(q_{i}\right)\mathop{}\!\mathrm{d}i\right).$ (14)
Here, the magnitude of the liquidity fee drives the trade-off between the participation of liquidity providers (LP) and trading costs. A lower fee $f$ results in fewer LPs offering liquidity, leading to increased unrealized gains. In contrast, a higher fee increases trading costs, potentially outweighing the benefits of increased LP participation.
The trade-off is illustrated in the left panel of Figure 4. The optimal fee $f^{\star}$ that minimizes the single-pool implementation shortfall is equal to $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)$, where $W(\cdot)$ represents the Lambert function.
@@
For any single-pool fee $f\geq 0$, there exists a set of fees $\{h,\ell\}$ for a two-pool fragmented market, where $h=f$ and $h>\ell$, that guarantees an equal or reduced implementation shortfall in a fragmented market compared to the single-pool market.
Furthermore, when $f>\frac{\Gamma}{Q}$, the fee structure $\{h,\ell\}$ can be chosen to ensure a strictly lower implementation shortfall in the fragmented market.
Proposition 3 suggests that fragmentation with multiple fee levels improves market quality. Specifically, it is always possible to devise a fee structure in a fragmented market that yields a (weakly) lower implementation shortfall than a single-fee market. The logic is as follows: First, the highest fee in the fragmented market is set equal to the single pool fee, ensuring that the marginal LP participating the market is the same across both scenarios (i.e., the LP with endowment $\underline{q}=\frac{\Gamma}{f}$). This condition guarantees the same level of realized gains from trade in fragmented and non-fragmented markets. Second, a lower fee is then chosen for another pool to attract liquidity providers with higher token endowments, resulting in lower trading costs. This combination of reduced trading costs and unchanged gains from trade leads to a lower implementation shortfall in a fragmented market.
Figure 4: Implementation shortfall and market structure
The left panel plots the implementation shortfall for large liquidity traders on a single pool, as a function of the pool fee $f$. The right panel plots the implementation shortfall for large liquidity traders on both single and fragmented pools as a function of the gas cost $\Gamma$. The single pool fee is set to the optimal level $f^{\star}$, whereas the fragmented pool fee menu is set to $\left(h,\ell\right)=\left(f^{\star},\frac{1}{2}f^{\star}\right)$.
Corollary 1.
There exists a fee menu in a two-pool fragmented market structure that achieves an equal or lower implementation shortfall than on any single-fee pool configuration. Further, the fee menu can be chosen to achieve strictly lower implementation shortfall in fragmented markets if $f^{\star}=gW^{-1}\left({\rm e}\frac{gQ}{\Gamma}\right)>\frac{\Gamma}{Q}$.
Corollary 1 emerges as a particular case of Proposition 3, with the assumption that the single-pool operates at its optimal fee level. In essence, if a fragmented fee structure can be designed to achieve a lower implementation shortfall compared to an arbitrary single-pool fee, then a fee structure that dominates the optimally set single-pool fee achieves a lower implementation shortfall than any single-fee pool.
The right panel of Figure 4 illustrates the scenario where the lower fee in a fragmented market is set to half the optimal fee of a single pool, denoted as $\ell=\frac{1}{2}f^{\star}$. When gas prices are sufficiently low, LPs with large endowments choose to provide liquidity on the low-fee pool, leading to reduced trading costs and, consequently, a lower implementation shortfall. As gas prices increase, the implementation shortfall rises in both single-pool and fragmented markets, primarily because more LPs are priced out which results in higher unrealized gains. Additionally, the shortfall increases more rapidly with rising gas prices in the fragmented market as LPs switch over from the low- to the high-fee pool — that is, the result in Proposition 2. At very high gas prices, all LPs converge in the high-fee pool, effectively collapsing the fragmented market into a single pool with the optimal fee level.
@@ 2.3 Model implications and empirical predictions
Prediction 1 follows directly from Proposition 2 and Figure 3. A higher gas price increases the fixed cost of active liquidity management, particularly so for smaller liquidity providers. In response, LPs with lower endowments migrate to the high-fee pool where they trade less often.
@@
Prediction 2 follows from the equilibrium discussion in Section 2.1. Liquidity providers with large token endowments ($q_{i}>q_{\text{t}}$) deposit them in the low-fee pool since they are better positioned to actively manage liquidity due to economies of scale. LPs with lower endowments ($q_{i}\leq q_{\text{t}}$) either stay out of the market or choose pool $H$ which allows them to offer liquidity in a more passive manner.
@@
This figure illustrates equilibrium liquidity supply on a fragmented market. We simulate 100,000 liquidity endowments from the truncated Pareto distribution in (1). The top panel illustrates the simulated distributions: LPs to the left (right) of the marginal trader $q_{\text{t}}^{\star}$ provide liquidity on pool $H$ (pool $L$, respectively). The bottom left panel shows the gap in average liquidity deposits for LPs participating on the two pools. The bottom right panel shows the gap in the number of LPs participating on the two pools. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
Figure 5 illustrates this prediction through a Monte Carlo simulation. We plot the equilibrium liquidity supply decisions of 100,000 LPs with endowments drawn from density (1) and $Q=3$. The top panel highlights three groups of liquidity providers: low-endowment LPs (in green) that are being rationed out of the market due to high gas cost, medium-endowment LPs (blue) that deposit liquidity on pool $H$, and high-endowment LPs (orange) that choose the low-fee pool $L$.
@@
Next, Prediction 3 deals with differences between incoming trades on the two liquidity pools. For a wide range of model parameters, incoming order flow on the low fee pool consists of a large number of small trades, and occasional large trades. In contrast, there are few trades on the high fee pool, but they are all relatively large. The model can therefore reconcile two apparently conflicting patterns: one liquidity pool captures most of the trading volume, while the largest trades are executed on the competitor. Figure 6 illustrates the prediction through a Monte Carlo simulation of the model.
Figure 6: Trade volume and trade size on fragmented markets
The left panel illustrates the average trading volume per unit of time in both pools, using 100,000 Monte Carlo simulations with a discrete time increment of $\Delta t=0.1$. The right panel illustrates the average trade size on a given liquidity pool, conditional on having a non-zero volume in the discrete time interval $\Delta t$. If a large liquidity trader arrives, it depletes both pools – i.e., she trades amounts $\mathcal{L}_{\ell}$ and $\mathcal{L}_{h}$ on the low- and high-fee pools, respectively. Parameter values: $h=1$, $\ell=0.75$, $\lambda=0.5$, $\theta=0.66$, $Q=3$, and $\Gamma=1$.
@@
The bottom left panel highlights the clientele effect: that is, the average deposit is higher on pool $L$. Due to the skew of the Pareto distribution, however, there are more LP accounts active on pool $H$ than on pool $L$ for a wide range of parameter values.
Prediction 5:
LPs update liquidity more frequently on the low-fee than on the high-fee pool.
Prediction 5 is a consequence of Lemma 1. Liquidity cycles on pool $L$ are shorter than on pool $H$, since LPs on the low-fee pool trade against both small and large orders, rather than only against large orders on the high-fee pool. Consequently, we expect LPs on pool $L$ to actively manage their liquidity positions.
@@
A larger gas price leads to more frequent liquidity updates on the low-fee pool.
An increase in gas price leads to some LPs switching from the low- to the high-fee pool. As a result, liquidity supply on the low-fee pool drops, leading to a shorter cycle as incoming order flow depletes the pool at a faster pace.
@@ 3.1 Sample construction
To enhance our dataset, we combine the Subgraph data with public Ethereum data available on Google Big Query to obtain the position of each transaction in its block, as well as the gas price limit set by the trader and the amount of gas used.
@@ 3.2 Liquidity fragmentation patterns
This table reports descriptive statistics for variables used in the empirical analysis. Pool size is defined as the total value locked in the pool’s smart contract at the end of each day. We compute the balance on day $t$ as follows: we take the balance at day $t-1$ and add (subtract) liquidity deposits (withdrawals) on day $t$, as well as accounting for token balance changes due to trades. The liquidity balance on the first day of the pool is taken to be zero. End of day balances are finally converted to US dollars. Daily volume is computed as the sum of US dollar volume for all trades in a given pool and day. Liquidity share (Volume share) is computed as the ratio between a pool size (trading volume) for a given fee level and the aggregate size of all pools (trading volumes) for the same pair in a given day. Trade size and Mint size are the median trade and liquidity deposit size on a given pool and day, denominated in US dollars. Trade count represents the number of trades in a given pool and day. LP wallets counts the unique number of wallet addresses interacting with a given pool in a day. The liquidity yield is computed as the ratio between the daily trading volume and end-of-day TVL, multiplied by the fee tier. The price range for every mint is computed as the difference between the top and bottom of the range, normalized by the range midpoint – a measure that naturally lies between zero and two. The impermanent loss is computed as in Heimbach, Schertenleib, and Wattenhofer (2022) for a position in the range of 95% to 105% of the current pool price, with a forward-looking horizon of one hour. Finally, mint-to-burn and burn-to-mint times are defined as the time between a mint (burn) and a subsequent burn (mint) by the same address in the same pool, measured in hours. Mint-to-burn and burn-to-mint are recorded on the day of the final interaction with the pool.
@@
One concern with measuring average mint size is just-in-time liquidity provision (JIT). JIT liquidity providers submit very large and short-lived deposits to the pool to dilute competitors on an incoming large trade; they immediately withdraw the balance in the same block after executing the trade. In our sample, JIT liquidity provision is not economically significant, accounting for less than 1% of aggregate trading volume. However, it has the potential to skew mint sizes to the right, particularly in low-fee pools, without providing liquidity to the market at large. We address this issue by (i) filtering out JIT mints using the algorithm in Appendix D and (ii) taking the median liquidity mint size at day-pool level rather than the mean.
@@
The empirical pattern in Figure 8 echoes liquidity cycles as described in Section 2. Liquidity deposit tokens in Uniswap pools and then patiently wait for days until incoming order flow exhausts their position (i.e., posted liquidity no longer earns fees). Once this happens, LPs quickly re-balance their position in a matter of minutes – by removing stale liquidity and adding a new position around the current price. The cycle is longer on high-fee pools for which trading volume is lower and liquidity takes longer to deplete.
Importantly, LPs do not seem to “race” to update liquidity upon information arrival as in Budish, Cramton, and Shim (2015). First, they very rarely manage their position intraday. Second, LPs on Uniswap typically do not remove in-range liquidity that stands to trade first against incoming order flow and therefore bears the highest adverse selection risk. Our results are consistent with Capponi and Jia (2021) who theoretically argue that LPs have low incentives to manage liquidity on news arrival, as well as with Capponi, Jia, and Yu (2022) who find no evidence of traders racing to trade on information on Uniswap v2.
@@
Each interaction with smart contracts on the Ethereum blockchain requires computational resources, measured in units of “gas.” Upon submitting a mint or burn transaction to the decentralized exchange, each liquidity provider specifies their willingness to pay per unit of gas, that is they bid a “gas price.” Traders are likely to bid higher prices for more complex transactions or if they require a faster execution. To generate a conservative daily benchmark for the gas price, we compute the average of the lowest 1000 user gas bids for mint and burn interactions on day $t$, across all liquidity pools in the benchmark sample. These transactions are more likely to be plain vanilla deposits or withdrawals of liquidity, capturing the cost of a simple interaction with the decentralized exchange smart contract.

### v7

@@ 2.2 Pool fragmentation and market quality
We measure market quality by the realized gains from trade of liquidity traders. If the asset is traded on a sequence of pools, where $f_{k}$ and $T_{k}$ represent the fees and liquidity deposits on pool $k$, respectively, the expected gains from trade for liquidity traders are
$\text{GainsFromTrade}\left(\left\{f_{k}\right\}_{k}\right)=v\mathbb{E}\left[\sum_{k}\tau^{\star}\left(f_{k},\delta\right)\times\delta\right],$ (18)
where $\tau^{\star}\left(\delta\right)=T_{k}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f_{k}}{1+\delta}}\right\}\right\}$ is the optimal LT trade size, as defined in equation (7).
Suppose an asset is traded on a single pool that imposes a liquidity fee $f$. From equation (18) it follows that the gains from trade for an LT with private value $1+\delta$ are equal to
$\text{GainsFromTrade}\left(f\mid\delta\right)=v\delta T_{f}\min\left\{1,\frac{1+r}{r}\max\left\{0,1-\sqrt{\frac{1+f}{1+\delta}}\right\}\right\}.$ (19)
The total token supply on the single pool equals $T_{f}=e^{-\underline{q}_{f}\lambda}\left(\underline{q}_{f}+\lambda\right)$, where $\underline{q}_{f}$ is the marginal liquidity provider such that all LPs with endowment $q_{i}>\underline{q}_{f}$ join the pool. Here, the magnitude of the liquidity fee drives the trade-off between the participation of liquidity providers (LP) and trading costs. A lower fee $f$ results in fewer LPs offering liquidity, a lower token supply $T_{f}$, which limits gains from trade for liquidity traders. In contrast, a higher fee increases trading costs, potentially outweighing the benefits of increased LP participation.
@@
For any single-pool fee $f\geq 0$, there exists a set of fees $\{h,\ell\}$ for a two-pool fragmented market, where $h=f$ and $h>\ell$, that guarantees equal or higher gains from trade in a fragmented market compared to the single-pool market.
Proposition 3 suggests that fragmentation with multiple fee levels improves market quality. Specifically, it is always possible to devise a fee structure in a fragmented market that yields (weakly) higher gains from trade than a single-fee market. The logic is as follows: First, the highest fee in the fragmented market is set equal to the single pool fee, ensuring that the marginal LP participating the market is the same across both scenarios (i.e., the LP with endowment $\underline{q}_{h}$). This condition guarantees the same aggregate liquidity supply in fragmented and non-fragmented markets. Second, a lower fee is then chosen for another pool to attract liquidity providers with higher token endowments, resulting in larger trade sizes per unit of supplied liquidity. This combination of larger liquidity trades and unchanged aggregate liquidity supply leads to higher gains from trade in a fragmented market.
Figure 5: Gains from trade and market structure
The figure plots the expected gains from trade across LTs,
$\text{GainsFromTrade}=\int_{0}^{\Delta^{2}-1}v\delta\tau^{\star}\left(\delta\right)\phi\left(\delta\right)\mathop{}\!\mathrm{d}\delta,$
on both a single pool with a high fee as well as on fragmented pools, as a function of the gas cost $\Gamma$. Parameter values: $r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
Figure 5 illustrates the result. As gas price increase, the gains from trade drop in both single-pool and fragmented markets, primarily because more LPs are priced out which results in lower liquidity supply and higher price impact. Nevertheless, irrespective of the level of gas costs, the gains from trade are higher in the fragmented market.
We note that the argument discussed in this section is valid for any single-fee pool, including an optimally designed one. In essence, if a fragmented fee structure can be designed to achieve higher gains from trade compared to an arbitrary single-pool fee, then a fee structure that dominates the optimally set single-pool fee achieves higher gains from trade than any single-fee pool.
@@ 2.3 Model implications and empirical predictions
Prediction 1 follows directly from Proposition 2 and Figure 4. A higher gas price increases the cost of liquidity re-balancing. Given that re-balancing is more frequently required in the low-fee pool due to more intense arbitrage activity, liquidity providers, particularly those with smaller endowments, optimally migrate to the high-fee pool in response to a gas cost increase.
@@
Prediction 2 follows from the equilibrium discussion in Proposition 1. Liquidity providers with large token endowments ($q_{i}>q_{\text{t}}$) deposit them in the low-fee pool since they are better positioned to actively manage liquidity due to economies of scale. LPs with lower endowments ($q_{i}\leq q_{\text{t}}$) either stay out of the market or choose pool $H$ which allows them to offer liquidity in a more passive manner. Figure 6 illustrates this prediction by overlaying optimal pool choices on the distribution of LP endowments. Low-endowment LPs (in blue) that are being rationed out of the market due to high gas cost, medium-endowment LPs (gray) that deposit liquidity on pool $H$, and high-endowment LPs (red) that choose the low-fee pool $L$.
@@
This figure illustrates the endowment distribution of LPs and their choice of pools in a fragmented market. First, liquidity providers to the left of $\underline{q}_{h}$ do not provide liquidity on either pool. Next, LPs to the left (right) of the marginal trader $q_{\text{t}}^{\star}$ provide liquidity on pool $H$ (pool $L$, respectively). $r=0.001$, $h=2$, $\ell=1$, $\lambda=1$, $\eta=0.1$, $\Gamma=20$, and $\Delta=1.1\left(1+r\right)\sqrt{1+h}$.
@@
Next, Prediction 3 deals with differences between incoming trades on the two liquidity pools. If liquidity traders and arbitrageurs find it optimal to trade on pool $H$ since $\delta>h$, then they would also trade on pool $L$ since $h>\ell$ and therefore $\delta>\ell$. However, the opposite is not true: LTs and arbitrageurs with $\delta\in\left[\ell,h\right)$ only trade on pool $L$. In equilibrium, only a fraction of traders with sufficiently high private values are drawn to pool $H$.
@@
In a fragmented market equilibrium, the liquidity yield is higher on the low fee pool than on the high fee pool.
Prediction 4 is a consequence of Proposition 1. The high fee pool offers better protection against adverse selection and re-balancing costs. If the low fee pool attracts a positive market share, then it necessarily compensates with a higher liquidity yield.
Prediction 5:
@@
LPs re-balance liquidity more frequently on the low-fee than on the high-fee pool.
Liquidity providers re-balance their positions in a pool charging a fee $f$ only when the magnitude of news exceeds a threshold, specifically if $\delta>(1+f)(1+r)^{2}-1$. The likelihood of re-balancing given the news is $(1-\frac{\sqrt{1+f}(1+r)}{\Delta})$. Consequently, the duration of a liquidity cycle, expressed as $\frac{1}{\eta(1-\frac{\sqrt{1+f}(1+r)}{\Delta})}$, increases in the pool fee level.
Prediction 7:
Adverse selection cost is higher on the low fee pool than on the high fee pool.
This prediction follows directly from Lemma 2: a higher pool fee serves as a deterrent to arbitrageurs, particularly if the size of news remains below a threshold.
@@ 3.1 Sample construction
To enhance our dataset, we combine the Subgraph data with public Ethereum data available on Google Big Query to obtain the position of each transaction in its block, as well as the gas price limit set by the trader and the amount of gas used. Finally, we obtain block-by-block liquidity snapshot data across multiple price ranges from Kaiko.
@@ 3.2 Liquidity fragmentation patterns
This table reports descriptive statistics for variables used in the empirical analysis. Pool size is defined as the total value locked in the pool’s smart contract at the end of each day. We compute the balance on day $t$ as follows: we take the balance at day $t-1$ and add (subtract) liquidity deposits (withdrawals) on day $t$, as well as accounting for token balance changes due to trades. The liquidity balance on the first day of the pool is taken to be zero. End of day balances are finally converted to US dollars. Daily volume is computed as the sum of US dollar volume for all trades in a given pool and day. Liquidity share (Volume share) is computed as the ratio between a pool size (trading volume) for a given fee level and the aggregate size of all pools (trading volumes) for the same pair in a given day. Trade size and Mint size are the median trade and liquidity deposit size on a given pool and day, denominated in US dollars. Trade count represents the number of trades in a given pool and day. LP wallets counts the unique number of wallet addresses interacting with a given pool in a day. The liquidity yield is computed as the ratio between the daily trading volume and end-of-day TVL, multiplied by the fee tier. The price range for every mint is computed as the difference between the top and bottom of the range, normalized by the range midpoint – a measure that naturally lies between zero and two. Loss-versus-rebalancing is computed as the permanent price impact of swaps with a one-hour horizon. The impermanent loss is computed as in Heimbach, Schertenleib, and Wattenhofer (2022) for a position in the range of 95% to 105% of the current pool price, with a forward-looking horizon of one hour. Finally, mint-to-burn and burn-to-mint times are defined as the time between a mint (burn) and a subsequent burn (mint) by the same address in the same pool, measured in hours. Mint-to-burn and burn-to-mint are recorded on the day of the final interaction with the pool.
@@
One concern with measuring average mint size is just-in-time liquidity provision (JIT). As discussed for example in Capponi, Jia, and Zhu (2024), JIT liquidity providers submit very large and short-lived deposits to the pool to dilute competitors on an incoming large trade; they immediately withdraw the balance in the same block after executing the trade. In our sample, JIT liquidity provision is not economically significant, accounting for less than 1% of aggregate trading volume. However, it has the potential to skew mint sizes to the right, particularly in low-fee pools, without providing liquidity to the market at large. We address this issue by (i) filtering out JIT mints using the algorithm in Appendix D and (ii) taking the median liquidity mint size at day-pool level rather than the mean.
@@
The empirical pattern in Figure 8 echoes the re-balancing cycles as described in Section 2. Liquidity providers deposit tokens in Uniswap pools to trade against uninformed order flow. They only re-balance when their position becomes out-of-range and no longer earns fees. Once this happens, LPs quickly adjust their position in a matter of minutes – by removing stale liquidity and adding a new position around the current price. The re-balancing cycle tends to be longer on high-fee pools, where arbitrageurs only move the price outside the range if the asset value innovation is large enough.
We note that LPs do not seem to “race” to update liquidity upon information arrival as in Budish, Cramton, and Shim (2015). First, they very rarely manage their position intraday. Second, LPs on Uniswap typically do not remove in-range liquidity that stands to trade first against incoming order flow and therefore bears the highest adverse selection risk. Our results are consistent with Capponi and Jia (2021) who theoretically argue that LPs have low incentives to compete with arbitrageurs on news arrival, as well as with Capponi, Jia, and Yu (2022) who find no evidence of traders racing to trade on information on Uniswap v2. In our model, LPs tend to re-balance their position after an arbitrageur executed their trade.
Next, we examine the behavior of liquidity takers (LT). According to our model, small orders are typically routed to low-fee pools, while larger orders are split between both low- and high-fee pools. Figure 9 provides empirical evidence supporting this claim. We use liquidity snapshot data from Kaiko on USDC-WETH pools to simulate the optimal routing strategy for trades of various sizes for the last block of each day in our sample. This simulation considers both the price impact of trades and the associated liquidity fees. In line with our model, we find that trades smaller than 150 ETH (approximately \$450,000) optimally route over 90% of their size to the low-fee pool. Conversely, larger trades distribute their volume more evenly, with up to 40% being executed in high-fee pools.
Figure 9: Optimal order routing on Uniswap v3 pools
This figure displays the optimal order split for purchasing ETH using USDC across various trade sizes, on USDC-WETH Uniswap v3 pools with liquidity fees of 5 and 30 basis points. Order execution is optimized to minimize trading costs, encompassing both price impact and liquidity fees. We use liquidity distribution snapshots data from Kaiko, and focus on the last Ethereum block of each day from May 4, 2021, to July 15, 2023.
@@
Each interaction with smart contracts on the Ethereum blockchain requires computational resources, measured in units of “gas.” Upon submitting a mint or burn transaction to the decentralized exchange, each liquidity provider specifies their willingness to pay per unit of gas, that is they bid a “gas price.” Traders are likely to bid higher prices for more complex transactions or if they require a faster execution. To generate a conservative daily benchmark for the gas price, we compute the average of the lowest 1000 user gas bids for mint and burn interactions on day $t$, across all liquidity pools in the benchmark sample.